# On the reconstruction limits of complex networks

### v2

@@ II.2.3 Optimal reconstruction performance
where $\bm{a}^{*}$ denotes the adjacency of the true graph $g^{*}$, such that $a_{ij}^{*}$ counts the number of edges connecting the nodes $i$ and $j$ (we use the convention that $a_{ii}^{*}$ is always a multiple of 2) and $\bm{\pi}(x)=[\pi_{ij}(x)]_{ij}$ is the predicted matrix of the posterior marginal probabilities of the edge occupancy for some model $M$. We show in Appendix IV.4 that, provided that the data is generated with $M^{*}$ and the reconstruction is performed with $M$, the posterior entropy and the expected posterior loss are equal if $M$ is equal to $M^{*}$ in distribution. Consequently, the reconstructability is linearly related to the posterior loss as follows:
$U(G^{*}|X^{*})=1-\frac{\mathbb{E}_{G^{*},X^{*}}\quantity[\mathcal{L}\Big{(}\bm{A}^{*},\bm{\pi}(X^{*})\Big{)}]}{H(G^{*})}\,,$ (22)
@@
Fig. 4: Comparison between reconstructability and different performance metrics: (a) posterior loss (Eq. (21)), (b) mean error $\binom{N}{2}^{-1}\sum_{i<j}|a_{ij}-\pi_{ij}(x)|$, (c) area under the receiver operating characteristic curve (AUC) and (d) Jaccard similarity (see Ref. (Peixoto, 2019, Eq. 11)). Each point shows a different realization of the Glauber dynamics whose graphs are generated from the Erdős-Rényi model with $N=100$ nodes and $E=250$ edges, and whose initial conditions are random. Reconstructions are performed with the same model, whose parameters are fixed to those used for generating the data. We used time series of $T=500$ time steps (as in Fig. 2, the parameters other than the coupling constant $J$ are specified in Table 2). We generated 24 realizations of the process for each value of $J$ and used 30 different coupling values uniformly spaced between 0 and 0.5. These coupling values are fixed during inference. The colors indicated in the legend show the value of $J$ associated with the point (only 6 colors are shown for conciseness). Finally, we show the determination coefficients $R^{2}$ relating the performance metrics to $U(G^{*}|X^{*})$ in each plot. For panel (a), we used Eq. (22) directly to evaluate the determination coefficient, and for panels (b) and (d), we used standard linear regression to find the slope and estimate $R^{2}$. For panel (c), because the scaling is not linear like the other cases, we used instead log-linear regression to estimate $R^{2}$.
Figure 4 shows further numerical evidence of the relationship between the reconstructability of the Glauber model and reconstruction performance measures, including the posterior loss. In Figs. 4(a) and (b), we show how $U(G^{*}|X^{*})$ is well correlated with metrics quantifying the recontruction error, such as the posterior loss and the mean error. Similarly, Figs. 4(c) and (d) show that the reconstructability is positively correlated with the area under the receiver operating characteristic curve (AUC) and the Jaccard similarity Peixoto (2019), both measuring the similarity between the true and reconstructed graphs.
II.2.4 Reconstructability of hierarchical Bayesian models
@@
Consider the case where a TDG model with variables $(G^{*},X^{*})$ also includes hyperparameters, denoted $\theta^{*}$ with probability distribution $P(\theta^{*})$, such that $G^{*}$ is conditioned on $\theta^{*}$—i.e., $P(\theta^{*},G^{*})=P(\theta^{*})P(G^{*}|\theta^{*})$. Let $\hat{\theta}$ and $\hat{G}$ be the reconstructed random parameters and graph, respectively, which are reconstructed from $X^{*}$ via some distribution $P(\hat{\theta},\hat{G}|X^{*})$. The random variables $(\theta^{*},G^{*})$ are related to those of the reconstruction model $(\hat{\theta},\hat{G})$ via $X^{*}$ as follows:
@@
where, again assuming that we use a Bayesian reconstruction model $M$, we let $P(\hat{\theta}=\vartheta,\hat{G}=g|X^{*}=x^{*})=P(\theta=\vartheta,G=g|X=x^{*})$, which is given by Eq. (23). In this case, the mutual information between $(\theta^{*},G^{*})$ and $(\hat{\theta},\hat{G})$ can be bounded using the following data processing inequality:
@@
II.3 Data-driven reconstructability and model selection
Until now, we have assumed that the TDG model $M^{*}$ was known to compute the mutual information $I(G^{*};X^{*})$. Outside of theoretical settings however, the TDG process is typically unknown. Hence, we generally cannot evaluate the true reconstruction limit, although we may have access to many realizations of $X^{*}$ which should help get closer to it. Three remarks are in order. First, the reconstructability is independent of the observations; it strictly depends on $M^{*}$. Second, any generative model $M$—unlike nongenerative ones—has a reconstructability, noted $U(G|X)$, which is calculated identically to Eq. (18). In other words, the condition that $M$ is capable of generating new data is crucial to our ability to calculate a reconstructability value. However and thirdly, the reconstructability $U(G|X)$ of $M$ differ in two ways from $U(G^{*}|X^{*})$ related to the actual reconstruction limit of the data: (i) their values are potentially different and (ii) the data generation process is $M^{*}$, not $M$. Consequently, we can leverage the reconstructability of $M$, with these considerations in mind, to get a data-driven proxy of the true upper bound $U(G^{*}|X^{*})$.
II.3.1 Reconstruction index based on information gain
@@
Note that the expectation of the information gain yields back the mutual information between $G$ and $X$, i.e., $\mathbb{E}_{X}\quantity[\mathcal{I}_{M}(X)]=I(G;X)$. The information gain measures the reduction in the entropy of a variable $G$ achieved by learning the state $x$ of another variable $X$. It is primarily used in feature selection, especially decision tree training, where it is used as a criterion for how to best split the data (Mitchell, T. M., 1997, Chapter 3). Like the mutual information, the information gain can be shown to be non-negative (see Appendix IV.5) and upper-bounded:
@@
Like the reconstructability, the reconstruction index $\Psi_{M}$ is bounded between 00 and $1$. However its interpretation is more subtle, as we will see in the following sections. Indeed, the information gain, on which the reconstruction index is based, is the Kullback-Leibler (KL) divergence between the posterior and the prior of the reconstruction model. As a result, it quantifies how different the posterior of the reconstruction model is from the prior. When $\Psi_{M}=0$, the posterior and the prior are identical—no information is gained from knowing the data. On the other hand, when $\Psi_{M}=1$, the posterior probability mass is entirely located on a single graph, which is reflected in the fact that the KL divergence is maximized.
II.3.2 Interpretation of the reconstruction index under incorrect assumptions
@@
Fig. 5: Effect of varying the coupling constant on the validity of the reconstruction index. We generated time series of the Glauber dynamics with fixed $J^{*}=0.3$ on Erdős-Rényi graphs with $N=100$ nodes and $E=250$ edges, then reconstructed the graphs using the same Glauber model with other coupling constants $J$, used during the inference. Panel (a) shows the relationship between the reconstruction index $\Psi_{M}$ and the posterior loss $\mathcal{L}(\bm{a}^{*},\bm{\pi})$ between the true graphs and the posterior—each point corresponding to a different realization of the TDG process (graph and observations) from which we reconstructed the graph. Panels (b–d) respectively show the reconstruction index $\Psi_{M}$, posterior loss $\mathcal{L}(\bm{a}^{*},\bm{\pi})$, and evidence cross-entropy $\mathcal{H}_{M,M^{*}}$ (Eq. (32)) as functions of $J$. The dashed vertical line shows where $J=J^{*}$. We color coded the points according to $J$, as shown in the legend, including the true value $J^{*}$ (grey squares). As in Fig. 4, we show the linear relationship between the reconstructability and the posterior loss (Eq. (22)) with the dashed line in (d). Glauber time series were generated with $T=500$ time steps, and we generated 24 realizations with random initial conditions for each value of $J$ between 0 and 0.8 (like in Fig. 4, we show only a few values in the legend of (d)). In panels (b–d), we show the 90% confident intervals around the mean (displayed by the markers), although they are too small to be visible.
@@
We must be careful in our interpretation of the reconstruction index, as its value can be misleading if not used in the correct way. Figure 5 shows how the reconstruction index behaves when the reconstruction model is incorrect to different extents. In this example, we generated time series of the Glauber dynamics with a given coupling constant $J^{*}$, and reconstructed the graphs using the same Glauber model, but typically with an erroneous coupling constant $J\neq J^{*}$. As we can see in Fig. 5(a), the reconstruction index keeps increasing as $J$ gets larger, even when it gets larger than $J^{*}$. While the reconstruction index is larger for $J>J^{*}$, the posterior loss actually shows, as expected, that the reconstruction becomes worse. Figure 5(a) also illustrates how the reconstruction index correlates with the posterior loss, depending on $J$. Indeed, as $J$ gets closer to $J^{*}$, the reconstruction index converges to the true reconstructability of the reconstruction model, which increasingly becomes linearly related to the posterior loss as previously shown (see the Appendix IV.4) In this regime, the reconstruction index is a good proxy of the true reconstructability because $M$ properly approximates the behavior of $M^{*}$.
@@
The behavior of the reconstruction index when the reconstruction model is incorrect raises some important remarks. Recall that, fundamentally, the reconstruction index is a normalized version of the KL divergence between the posterior and the prior of the reconstruction model. Therefore, it is perhaps not surprising that we lose the correspondence between reconstruction index and performance established in Sec. II.2.3 when the model is incorrect. Indeed, having a posterior that is very different from the prior implies a high reconstruction index even though the posterior distribution is actually wrong.
@@
Another perspective of this observation reveals that maximizing the reconstruction index can lead to inadequate modeling of the observed data. Consider the following alternative but equivalent form of the information gain:
@@
In this formulation, the two terms—i.e., the expected log-predictive and the log-evidence, respectively—are in opposition. The former term is maximized when the model is good at describing the data using the posterior graphs, while the latter is maximized when the model describes the data correctly altogether. The log-evidence is even used as a measure of goodness-of-fit for model selection, as we will see in the next section. Yet, maximizing the information gain is equivalent to maximizing the expected log-predictive and minimizing the log-evidence, which is why incorrect models may be selected by this criterion. Following these remarks, we devote the next section to describing a principled approach to adequately interpret the reconstruction index and use it in the context of data-driven reconstruction.
II.3.3 Role of evidence-based model selection
@@
$B_{M_{1},M_{2}}(x)=\frac{P(X_{1}=x)}{P(X_{2}=x)}\,,$ (31)
where $P(X_{i}=x)$ is the evidence of model $M_{i}$ for $x$. If $B_{M_{1},M_{2}}(x)>1$, $M_{1}$ is better supported by the data $x$ than $M_{2}$. From an information-theoretic perspective, the minimization of the evidence CE—which is equivalent to maximizing the evidence—can be shown to be a necessary condition for finding the TDG model. Indeed, for a reconstruction model $M$ with evidence distribution $P(X)$, the evidence CE is expressed as
$\begin{split}\mathcal{H}_{M^{*},M}&=-\mathbb{E}_{X^{*}}\quantity[\log P(X=X^{*})]\,,\\ &=-\sum_{x^{*}\in\mathcal{X}}P(X^{*}=x^{*})\log P(X=x^{*})\,.\end{split}$ (32)
@@
The problem of evidence-based model selection is the computation of the evidence itself which is often intractable in practice. In fact, this is the case for most graph models, where the evaluation of the evidence requires graph enumeration. This problem also arises in the evaluation of both the information gain and the mutual information. Fortunately, the same numerical techniques can be used to evaluate the evidence and the information gain simultaneously, as the two are related to each other. In Ref. Murphy et al. (2024), we showed that variational mean-field methods provide efficient approximations for both the mutual information and the evidence. The same techniques are used here (see Sec. IV.6).
Model selection is crucial for the validity of the reconstruction index as a proxy of reconstruction performance. Suppose we observed many observations $(x_{1},x_{2},...)$ and used the reconstruction model $M^{*}$. The empirical average of the information gain of some model $M$ converges to $\mathbb{E}_{X^{*}}\quantity[\mathcal{I}_{M}(X^{*})]$. Now, assume for a moment that $M$ in fact maximizes the expected log-evidence. This implies, as we mentioned before, that $X^{*}$ and $X$ are equal in distribution. In turn, the empirical information gain becomes equal to the mutual information for model $M$, i.e., $I(X;G)$, which we recall is the reconstruction limit of $M$. This means that empirical information gain converges to the reconstructability and that ultimately the reconstruction index naturally extends the concept of reconstruction limit to real systems observed only through data.
@@
II.4 Network reconstructability in empirical networks
@@
Fig. 6: Reconstruction from spontaneous neuronal activity in the mouse brain Steinmetz et al. (2019); Stringer et al. (2019): (a) Raster plot of the 1462 monitored neurons, (b) reconstruction of the probe network using different reconstruction models and (c) reconstructability diagram. In panel (a), the neurons are ordered by the probe they were measured from. Each spike is represented in blue. Panel (b) shows the posterior average network projected onto the probes, as predicted by each reconstruction model where rows correspond to different graph models (see Appendix IV.1), and columns to different dynamics models (see Table 2). The color of an edge connecting two probes shows the absolute number of edges and the thickness indicates the average proportion among all the edges. The size of the probe nodes is proportional to the number of neurons monitored by the probe, and the color indicates the measured number of spikes. The node locations correspond to the actual probe locations in the mouse brain obtained from Steinmetz et al. (2019). The reconstruction index as a function of the model log evidence is shown in panel (c), comparing the different models. Small markers are estimated by a single Markov chain and large markers are the average of these estimations. In these experiments, the parameters of the graph prior and and likelihood are inferred jointly with the graph. For additional details about the inference procedure, we refer to Appendix IV.9.
Reconstructing empirical graphs represents a technical and conceptual challenge. The true network structure being unknown, it is hard to quantify how close the predicted graph is to the true one, let alone calculate its actual reconstructability. As explained in the previous section, the reconstruction index offers a means to predict the reconstruction limit without relying on the true network structure, provided that the reconstruction model is appropriately selected. Building on this insight, we demonstrate how our framework can be applied to empirical systems.
II.4.1 Reconstruction from empirical neuronal spiking data

### v3

@@ III.3 Optimal reconstruction performance
where $\bm{a}^{*}$ denotes the adjacency of the true graph $g^{*}$, such that $a_{ij}^{*}$ counts the number of edges connecting the nodes $i$ and $j$ (we use the convention that $a_{ii}^{*}$ is always a multiple of 2) and $\bm{\pi}(x)=[\pi_{ij}(x)]_{ij}$ is the predicted matrix of the posterior marginal probabilities of the edge occupancy for some model $M$. We show in Appendix D that, provided that the data is generated with $M^{*}$ and the reconstruction is performed with $M$, the posterior entropy and the expected posterior loss are equal if $M$ is equal to $M^{*}$ in distribution. Consequently, the reconstructability is linearly related to the posterior loss as follows:
$\Psi^{*}\approx 1-\frac{\mathbb{E}_{G^{*},X^{*}}\quantity[\mathcal{L}\Big{(}\bm{A}^{*},\bm{\pi}(X^{*})\Big{)}]}{H(G^{*})}\,$, (22)
@@
Fig. 4: Comparison between reconstructability and different performance metrics: (a) posterior loss (Eq. (21)), (b) mean error $\binom{N}{2}^{-1}\sum_{i<j}|a_{ij}-\pi_{ij}(x)|$, (c) area under the receiver operating characteristic curve (AUC) and (d) Jaccard similarity (see Ref. (Peixoto, 2019, Eq. 11)). Each point shows a different realization of the Glauber dynamics whose graphs are generated from the Erdős-Rényi model with $N=100$ nodes and $E=250$ edges, and whose initial conditions are random. Reconstructions are performed with the same model, whose parameters are fixed to those used for generating the data. We used time series of $T=500$ time steps (as in Fig. 2, the parameters other than the coupling constant $J$ are specified in Table 2). We generated 24 realizations of the process for each value of $J$ and used 30 different coupling values uniformly spaced between 0 and 0.5. These coupling values are fixed during inference. The colors indicated in the legend show the value of $J$ associated with the point (only 6 colors are shown for conciseness). Finally, we show the determination coefficients $R^{2}$ relating the performance metrics to $\Psi^{*}$ in each plot. For panel (a), we used Eq. (22) directly to evaluate the determination coefficient, and for panels (b) and (d), we used standard linear regression to find the slope and estimate $R^{2}$. For panel (c), because the scaling is not linear like the other cases, we used instead log-linear regression to estimate $R^{2}$.
Figure 4 shows further numerical evidence of the relationship between the reconstructability of the Glauber model and reconstruction performance measures, including the posterior loss. In Figs. 4(a) and (b), we show how $\Psi^{*}$ is well correlated with metrics quantifying the recontruction error, such as the posterior loss and the mean error. Similarly, Figs. 4(c) and (d) show that the reconstructability is positively correlated with the area under the receiver operating characteristic curve (AUC) and the Jaccard similarity Peixoto (2019), both measuring the similarity between the true and reconstructed graphs.
III.4 Reconstructability of hierarchical Bayesian models
@@
[…]
@@
[…]
@@
IV Data-driven reconstructability and model selection
Until now, we have assumed that the TDG model $M^{*}$ was known to compute the mutual information $I(G^{*};X^{*})$. Outside of theoretical settings however, the TDG process is typically unknown. Hence, we generally cannot evaluate the true reconstruction limit, although we may have access to many realizations of $X^{*}$ which should help get closer to it. Three remarks are in order. First, the reconstructability $\Psi^{*}$ is independent of the observations; it strictly depends on $M^{*}$. Second, any generative model $M$ has a reconstructability, which is calculated identically to Eq. (18). In other words, the condition that $M$ is capable of generating new data is crucial to our ability to calculate a reconstructability value. However and thirdly, the reconstructability of $M$ differ in two ways from $\Psi^{*}$ related to the actual reconstruction limit of the data: (i) their values are potentially different and (ii) the data generation process is $M^{*}$, not $M$. Consequently, we can leverage the reconstructability of $M$, with these considerations in mind, to get a data-driven proxy of the true upper bound $\Psi^{*}$.
IV.1 Reconstruction index based on information gain
@@
Note that the expectation of the information gain yields back the mutual information between $G$ and $X$, i.e., $\mathbb{E}_{X}\quantity[\mathcal{I}_{M}(X)]=I(G;X)$. The information gain measures the reduction in the entropy of a variable $G$ achieved by learning the state $x$ of another variable $X$. It is primarily used in feature selection, especially decision tree training, where it is used as a criterion for how to best split the data (Mitchell, T. M., 1997, Chapter 3). Like the mutual information, the information gain can be shown to be non-negative (see Appendix E) and upper-bounded:
@@
Like the reconstructability, the reconstruction index $\psi_{M}$ is bounded between 0 and $1$. However, it differs mainly in that $\psi_{M}$ may yield different values for different datasets $x$. In addition, there is a subtle difference in their interpretations, as we will see in the following sections. Indeed, the information gain, on which the reconstruction index is based, is the Kullback-Leibler (KL) divergence between the posterior and the prior of the reconstruction model. As a result, it quantifies how different the posterior of the reconstruction model is from the prior. When $\psi_{M}=0$, the posterior and the prior are identical—no information is gained from knowing the data. On the other hand, when $\psi_{M}=1$, the posterior probability mass is entirely located on a single graph, which is reflected in the fact that the KL divergence is maximized.
IV.2 Interpretation of the reconstruction index under incorrect assumptions
@@
Fig. 5: Effect of varying the coupling constant on the validity of the reconstruction index. We generated time series of the Glauber dynamics with fixed $J^{*}=0.3$ on Erdős-Rényi graphs with $N=100$ nodes and $E=250$ edges, then reconstructed the graphs using the same Glauber model with other coupling constants $J$, used during the inference. Panel (a) shows the relationship between the reconstruction index $\psi_{M}$ and the posterior loss $\mathcal{L}(\bm{a}^{*},\bm{\pi})$ between the true graphs and the posterior—each point corresponding to a different realization of the TDG process (graph and observations) from which we reconstructed the graph. Panels (b–d) respectively show the reconstruction index $\psi_{M}$, posterior loss $\mathcal{L}(\bm{a}^{*},\bm{\pi})$, and evidence cross-entropy $\mathcal{H}_{M,M^{*}}$ (Eq. (32)) as functions of $J$. The dashed vertical line shows where $J=J^{*}$. We color-coded the points according to $J$, as shown in the legend, including the true value $J^{*}$ (grey squares). As in Fig. 4, we show the linear relationship between the reconstructability and the posterior loss (Eq. (22)) with the dashed line in (d). Glauber time series were generated with $T=500$ time steps, and we generated 24 realizations with random initial conditions for each value of $J$ between 0 and 0.8 (like in Fig. 4, we show only a few values in the legend of (d)). In panels (b–d), we show the 90% confident intervals around the mean (displayed by the markers), although they are too small to be visible.
@@
We must be careful in our interpretation of the reconstruction index, as its value can be misleading if not used in the correct way. Figure 5 shows how the reconstruction index behaves when the reconstruction model is incorrect to different extents. In this example, we generated time series of the Glauber dynamics with a given coupling constant $J^{*}$, and reconstructed the graphs using the same Glauber model, but typically with an erroneous coupling constant $J\neq J^{*}$. As we can see in Fig. 5(a), the reconstruction index keeps increasing as $J$ gets larger, even when it gets larger than $J^{*}$. While the reconstruction index is larger for $J>J^{*}$, the posterior loss actually shows, as expected, that the reconstruction becomes worse. Figure 5(a) also illustrates how the reconstruction index correlates with the posterior loss, depending on $J$. Indeed, as $J$ gets closer to $J^{*}$, the reconstruction index converges to the true reconstructability of the reconstruction model, which increasingly becomes linearly related to the posterior loss as previously shown (see the Appendix D) In this regime, the reconstruction index is a good proxy of the true reconstructability because $M$ properly approximates the behavior of $M^{*}$.
@@
The behavior of the reconstruction index when the reconstruction model is incorrect raises some important remarks. Recall that, fundamentally, the reconstruction index is a normalized version of the KL divergence between the posterior and the prior of the reconstruction model. Therefore, it is perhaps not surprising that we lose the correspondence between reconstruction index and performance established in Sec. III.3 when the model is incorrect. Indeed, having a posterior that is very different from the prior implies a high reconstruction index even though the posterior distribution is actually wrong.
@@
Maximizing the reconstruction index can also lead to inadequate modeling of the observed data. Consider the following alternative but equivalent form of the information gain:
@@
In this formulation, the two terms—i.e., the expected log-predictive and the log-evidence, respectively—are in opposition. The first term is maximized when the model is good at describing the data using the posterior graphs, while the second is maximized when the model describes the data correctly altogether. The log-evidence is even used as a measure of goodness-of-fit for model selection, as we will see in the next section. Yet, maximizing the information gain is equivalent to maximizing the expected log-predictive and minimizing the log-evidence, which is why incorrect models may be selected by this criterion. Following these remarks, we devote the next section to describing a principled approach to adequately interpret the reconstruction index and use it in the context of data-driven reconstruction.
IV.3 Role of evidence-based model selection
@@
$B_{M_{1},M_{2}}(x)=\frac{\zeta_{1}(x)}{\zeta_{2}(x)}\,$, (31)
where $\zeta_{M_{i}}(x)=P(X_{i}=x)$ is the evidence of model $M_{i}$ for $x$. If $B_{M_{1},M_{2}}(x)>1$, $M_{1}$ is better supported by the data $x$ than $M_{2}$. From an information-theoretic perspective, the minimization of the evidence cross-entropy (CE)—which is equivalent to maximizing the evidence—can be shown to be a necessary condition for finding the TDG model, whose evidence function is $\zeta_{M^{*}}(x)$. Indeed, for a reconstruction model $M$ with evidence function $\zeta_{M}(x)$, the evidence CE is expressed as
$\begin{split}\mathcal{H}_{M^{*},M}&=-\mathbb{E}_{X^{*}}\quantity[\log\zeta_{M}(X^{*})]\,,\\ &=-\sum_{x^{*}\in\mathcal{X}}\zeta_{M^{*}}(x^{*})\log\zeta_{M}(x^{*})\,.\end{split}$ (32)
@@
The problem of evidence-based model selection is the computation of the evidence itself which is often intractable in practice. In fact, this is the case for most graph models, where the evaluation of the evidence requires graph enumeration. This problem also arises in the evaluation of both the information gain and the mutual information. Fortunately, the same numerical techniques can be used to evaluate the evidence and the information gain simultaneously, as the two are related to each other. In Ref. Murphy et al. (2024), we showed that variational mean-field methods provide efficient approximations for both the mutual information and the evidence. The same techniques are used here (see Appendix F).
[…]
@@
V Network reconstructability in empirical networks
@@
Fig. 6: Reconstruction from spontaneous neuronal activity in the mouse brain Steinmetz et al. (2019); Stringer et al. (2019): (a) Raster plot of the 1462 monitored neurons, (b) reconstruction of the probe network using different reconstruction models and (c) reconstructability diagram. In panel (a), the neurons are ordered by the probe they were measured from. Each spike is represented in blue. Panel (b) shows the posterior average network projected onto the probes, as predicted by each reconstruction model where rows correspond to different graph models (see Appendix A), and columns to different dynamics models (see Table 2). The color of an edge connecting two probes shows the absolute number of edges and the thickness indicates the average proportion among all the edges. The size of the probe nodes is proportional to the number of neurons monitored by the probe, and the color indicates the measured number of spikes. The node locations correspond to the actual probe locations in the mouse brain obtained from Steinmetz et al. (2019). The reconstruction index as a function of the model log evidence is shown in panel (c), comparing the different models. Small markers are estimated by a single Markov chain and large markers are the average of these estimations. In these experiments, the parameters of the graph prior and likelihood are inferred jointly with the graph. For additional details about the inference procedure, we refer to Appendix I.
Reconstructing empirical graphs represents a technical and conceptual challenge. The true network structure being unknown, it is hard to quantify how close the predicted graph is to the true one, let alone calculate its actual reconstructability. In light of our exploration in Sec. IV, however, we have shown that the reconstruction index $\psi_{M}$ can offer a means to approximate the reconstructability, if certain conditions are met. Additionally, our analysis shows that a correctly calibrated reconstruction index $\psi_{M}$ predicts the error in a reconstructed network in comparison with the true one. Under these considerations, we present a principled method based on the reconstruction index to assess the validity of network reconstruction.
In this procedure, we assume $x$ to be some time series generated by a hidden process $M^{*}$, from which we wish to infer a network. Next, the procedure goes as follows:
  1. 1.
    Select a set of $d$ reconstruction model candidates $\mathcal{M}=\{M_{1},M_{2},...,M_{d}\}$.
  2. 2.
    For each candidate $M$, sample a reconstructed graph ensemble $\hat{\mathcal{G}}_{M}$ using the posterior of $M$.
  3. 3.
    Calculate for each candidate $M$ the evidence $\zeta_{M}(x)$ and the reconstruction index $\psi_{M}(x)$ using $\hat{\mathcal{G}}_{M}$ (see Sec. F for detail).
  4. 4.
    Choose the reconstruction index $\psi_{\hat{M}}(x)$ of the model $\hat{M}$ with the highest evidence:
    $\hat{M}=\operatorname*{arg\,max}_{M\in\mathcal{M}}\zeta_{M}(x)\,.$ (33)
    The index $\psi_{\hat{M}}(x)$ is the output of the procedure.
Some remarks about this method are in order. Indeed, as shown in Sec. IV.3, the validity of the reconstruction index heavily relies on the model’s aptitude to represent the data. Hence, it is paramount that an adequate set of model candidates is selected at first. These models may differ in their underlying assumptions, via their prior, hyper prior, and/or likelihood functions, as previously stated. If $\mathcal{M}$ contains a model that resembles the TDG process, this procedure will generate a reconstruction index $\psi_{\hat{M}}(x)$ that closely approximates the true reconstructability of the process $\Psi^{*}$.
Of course, determining if such a model is in $\mathcal{M}$ is hardly feasible experimentally. We work around this issue by performing a validation after the procedure via a predictive posterior check of $\hat{M}$. This validation proceeds as follows:
  1. 1.
    Generate a sample $\left\{\hat{x}_{1},...,\hat{x}_{K}\right\}$ of synthetic data with $\hat{M}$, assuming its parameters $(\theta,\phi)$ are sampled from the model’s posterior given $x$;
  2. 2.
    Calculate some test quantities $\tau(\hat{x}_{k})$, i.e., statistics used for comparison, for each sample $\hat{x}_{k}$, generating a set of samples $\mathcal{T}=\left\{\tau(\hat{x}_{1}),...,\tau(\hat{x}_{K})\right\}$;
  3. 3.
    Compare $\tau(x)$ with $\mathcal{T}$.
If $\tau(x)$ is typical in $\mathcal{T}$, then we can be sure $\hat{M}$ is statistically similar to $M^{*}$.
In what follows, we will demonstrate this method in two different empirical use cases. The first use case reflects a realistic modeling scenario, where only the time series on which reconstruction is performed are known. From a theoretical perspective, this use case presents challenges that falls outside of the scope of this paper (we will discuss those in the following section). The second use case lift those limitations by reconstructing real networks from synthetic time series.
V.1 Reconstruction from empirical neuronal spiking data